OpenStudy (anonymous):

PLEASE HELP!!! Convert -450° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

I got -7.8539 but what is that in pi form?

OpenStudy (anonymous):

just multiply -450 times \[\frac{ \pi }{ 180 }\] and reduce.

OpenStudy (anonymous):

\[-450\times \frac{ \pi }{ 180 }=-\frac{ 5\pi }{ 2 }\]

OpenStudy (anonymous):

That is your exact value. Then just use the decimal value for pi to get the approximation.

OpenStudy (anonymous):

thank you so much!!!!!!